OpenStudy (anonymous):

Hi, using factoring by grouping, solve 6x^3-6x^2-x+1

OpenStudy (bahrom7893):

okay let us observe what is similar here: 6x^3-6x^2-x+1, i can factor out 6x^2 from the first two terms: 6x^2(x-1) - x+1

OpenStudy (anonymous):

check that 1 is a solution to \[6x^3-6x^2-x+1=0\] and then divide by \[(x-1)\] to get \[(x-1)(6x^2-1)\]

OpenStudy (bahrom7893):

6x^2(x-1) - x+1 i can factor out the - sign from the next two terms: 6x^2(x-1) - (x-1), now I can factor out the whole x-1 (x-1) (6x^2-1)

OpenStudy (anonymous):

so zeros are \[1,-\frac{1}{\sqrt{6}}, \frac{1}{\sqrt{6}}\]

OpenStudy (bahrom7893):

x-1=0 6x^2-1=0 x=1 6x^2=1, x^2 = 1/6 => x1,2 = +or- sqrt(6)/6 and both me and satellite are right!